The ratio of men to women working for a company 4 to 5 . If there are 75 women working for the company, what is the total number of employees?

Answer by hamsanash1981@gmail.com(151)

Given that the ratio of men to woman is 4:5. Total ratio is 4+5=9
Let x be the total employees then,
Woman = x*5/9
75 = x*5/9
x = 75*9/5
The total employees are 135